# GaussRender: Learning 3D Occupancy with Gaussian Rendering

### v2

@@ Abstract
Understanding the 3D geometry and semantics of driving scenes is critical for safe autonomous driving. Recent advances in 3D occupancy prediction have improved scene representation but often suffer from spatial inconsistencies, leading to floating artifacts and poor surface localization. Existing voxel-wise losses (e.g., cross-entropy) fail to enforce geometric coherence. In this paper, we propose GaussRender, a module that improves 3D occupancy learning by enforcing projective consistency. Our key idea is to project both predicted and ground-truth 3D occupancy into 2D camera views, where we apply supervision. Our method penalizes 3D configurations that produce inconsistent 2D projections, thereby enforcing a more coherent 3D structure. To achieve this efficiently, we leverage differentiable rendering with Gaussian splatting. GaussRender seamlessly integrates with existing architectures while maintaining efficiency and requiring no inference-time modifications. Extensive evaluations on multiple benchmarks (SurroundOcc-nuScenes, Occ3D-nuScenes, SSCBench-KITTI360) demonstrate that GaussRender significantly improves geometric fidelity across various 3D occupancy models (TPVFormer, SurroundOcc, Symphonies), achieving state-of-the-art results, particularly on surface-sensitive metrics. The code is open-sourced at https://github.com/valeoai/GaussRender.
Figure 1: Comparison of rendered 3D predictions. Standard 3D Occupancy models trained on per-voxel losses result in physically implausible predictions (e.g., floating voxels, poorly localized surfaces, highlighted with orange ellipses) that maintain high 3D IoU but fail to produce consistent predictions. GaussRender enforces multi-view consistency, eliminating artifacts through learning projective constraints.
@@ 1 Introduction
The challenge in 3D occupancy prediction lies in achieving geometrically coherent reasoning from multi-view images. Existing methods [54, 15, 16, 30] typically optimize per-voxel predictions using standard 3D losses, such as cross-entropy, Dice [48], or Lovász [2]. However, these losses treat all voxels equally and fail to enforce spatial consistency between neighboring voxels, leading to fragmented or floating occupancy artifacts, as noticed in prior works [36] and illustrated in Fig. 1. Such artifacts (floating voxels, disjoint surfaces, and misaligned boundaries) often have a low impact on voxel-based segmentation losses, which give the same weight to all voxels. The consequences extend beyond mere visual artifacts: poor surface localization and unrealistic floating objects may significantly undermine downstream tasks such as free-space estimation or motion planning. The underlying issue is that conventional supervision methods lack mechanisms to penalize unrealistic spatial arrangements, a deficiency that becomes especially apparent when projecting 3D artifacts into 2D views. This observation motivates our key insight: integrating projective consistency into the training objective encourages the model to learn consistent and physically plausible geometries.
@@
We introduce GaussRender, a module that bridges the gap through differentiable rendering of 3D occupancy predictions. Our key idea is to enforce spatial consistency during training, by projecting both predicted and ground-truth voxels into 2D camera views using Gaussian splatting [23]. These projections serve as a supervision signal, allowing us to penalize 3D configurations that produce poor 2D projections, thereby enforcing a more coherent and geometrically plausible 3D structure. To achieve this, we apply two complementary supervision signals: (1) a semantic rendering loss that enforces local semantic coherence, and (2) a depth rendering loss that penalizes occlusion-disrupting artifacts. These rendering-based losses are applied alongside standard 3D supervision, such as cross-entropy, ensuring compatibility with existing occupancy learning frameworks. GaussRender enables rendering from arbitrary viewpoints. Its flexibility helps reduce occlusions, for instance, by leveraging elevated viewpoints that are less affected by horizontal obstructions from ground objects.
@@
We validate our approach on multiple datasets, including SurroundOcc-nuScenes [54], Occ3D-nuScenes [50], and SSCBench-KITTI360 [31]. Our experiments demonstrate that GaussRender significantly improves geometric fidelity across diverse architectures — ranging from multi-scale voxel-based models (SurroundOcc [54]), tri-plane models (TPVFormer [15]) to hybrid query-voxel models (Symphonies [19]). Across all settings and datasets, GaussRender consistently achieves state-of-the-art performance on classical metrics such as IoU and mIoU. Moreover, when evaluated using surface- and artifact-sensitive metrics, such as RayIoU [36], the improvements are even more pronounced. This is because the projective constraints enforced by our method promote surface continuity — ensuring that neighboring voxels agree when rendered from any viewpoint, thereby eliminating floating artifacts and discontinuities. Crucially, our approach is plug-and-play, integrating seamlessly with existing 3D occupancy frameworks without requiring any architectural modifications. Thanks to our efficient Gaussian rendering proxy, GaussRender incurs minimal computational and memory overhead compared to NeRF-based alternatives [17, 43]. Moreover, it introduces only a few hyperparameters, which are only loosely sensitive, ensuring straightforward integration into any pipeline.
@@
  • A rendering-based module that enforces semantic and geometric consistency in 3D occupancy prediction, eliminating spatial artifacts through projective constraints,
@@
  • A viewpoint selection strategy that amplifies supervision signals in geometrically complex regions.
  • State-of-the-art results on three standard benchmarks, with particular gains in surface-sensitive metrics.
@@ 2.1 Learning 3D semantic geometry from cameras
Discrete methods rely on regular 3D grids to stay as close as possible to the desired voxelic output. Bird’s-eye-view (BeV) projections [51, 27] significantly reduce memory cost but introduce a substantial bias in the intermediate representation due to height compression. Instead, tri-plane features [15] and tensor decompositions [62] map any 3D coordinate onto distinct planes or vectors performing interpolation to get distinct features. As opposed to BeV approaches, they do not suffer from height compression while still operating on a compressed representation. Octrees [39] preserves a full multi-scale 3D representation, coarse in uniform areas and fine where details are needed, resulting in a fast and memory-efficient representation [54]. Continuous architectures have emerged, using a set of Gaussians to represent the scene [16, 18] which are then discretized to obtain the voxelized output map. Furthermore, to better capture the overall instance semantics and scene context, query-based methods have been introduced [19]. They facilitate interactions between image and volume features using sparse instance queries.
@@
Figure 2: Overview of GaussRender. Our module enforces 3D-2D consistency via differentiable Gaussian rendering. First, both predicted and ground-truth voxel grids are ‘gaussianized’ by converting each voxel into a simple spherical Gaussian: the center $\mu$ is fixed at the voxel center, the scale $s$ is a simple fixed scaling of the original voxel dimensions, and features are directly transferred as the semantic ‘color’ $c$ — with only the opacity $o$ learned when voxel features are available. Next, virtual cameras are positioned in the scene (a fixed bird’s-eye view and a dynamic, arbitrarily placed camera as described in Sec. 3.2). The resulting 3D Gaussians are then projected into 2D using Gaussian splatting (Sec. 3.3), producing both semantic and depth renderings. These rendered views are compared against their ground-truth counterparts using an L1 penalty, ensuring enhanced spatial coherence and geometric consistency (Sec. 3.4).
@@
Complementary approaches have been proposed to enhance 3D occupancy predictions beyond architectural choices. Temporal aggregation uses past frames to resolve occlusions and refine geometric details [27, 59, 26, 46, 43], with extensions to 4D forecasting for dynamic scenes [24, 58, 53, 40]. Self-supervised methods reduces dependency on 3D annotations by generating pseudo-labels from monocular depth and segmentation [17, 7], but suffer from scale miscalibrations and label mismatches between 2D and 3D labels. Lastly, supervised rendering with lidar reprojections [43, 49] has been introduced to provide accurate distances. While more precise than pseudo-annotations, lidar reprojections face challenges such as signal sparsity, occlusions, and misalignment between the lidar and cameras. In particular, [43, 49] require temporal supervision from adjacent frames to compensate for lidar sparsity and occlusions. In contrast, our method achieves accurate supervision without these constraints, allowing more flexible and efficient 3D occupancy learning across different architectures.
@@ 3.1 Enforcing 3D-2D consistency
While the standard 3D losses (denoted as $L_{\text{3D}}$) ensure overall occupancy alignment with the ground truth, they treat each voxel independently and ignore spatial continuity. This limitation can lead to artifacts such as floating voxels or misaligned surfaces [36]. To overcome these issues, we introduce a 2D rendering loss, $L_{\text{2D}}$, which provides additional supervision by comparing rendered views of the predicted occupancy with corresponding ground-truth images. This loss enforces spatial coherence in the 2D projection, thereby penalizing 3D configurations that produce poor renderings. The overall training loss is defined as:
@@ 3.2 Camera Placement Strategy
GaussRender renders from arbitrary camera positions within the scene, rather than being restricted to the original sensor or adjacent frames [43, 17, 49]. This flexibility allows us to position virtual cameras anywhere in the 3D space. Fig. 3 illustrates how virtual cameras generate complementary constraints (both semantic and depth) to enhance voxel consistency.
The placement of these virtual cameras is crucial and depends on the reconstruction objective:
  • Visible Voxels: If the goal is to accurately reconstruct only the visible portions of the scene [50, 31], the rendering cameras are placed at the original sensor positions, as illustrated in Fig. 3 (a). This placement ensures that the inferred 3D structure remains consistent with the sensor’s view. In practice, for each batch, we randomly choose one of the provided camera positions.
  • Holistic 3D Reconstruction: For complete 3D reconstruction, including occluded regions [54], cameras are positioned more diversely. Exploring different viewpoints provides additional supervision for occluded areas, heavily penalizing aberrant configurations. In practice, we generate the virtual views with a simple rule: the camera is (1) elevated along $z$ axis with respect to its original position, (2) randomly translated in a close range to the ego-vehicle on the $xy$ plane. The resulting camera increases the field of view, looking at both visible and occluded parts of the scene (Fig. 3 (b)).
@@ 3.3 Gaussian rendering
For each camera, once its position is chosen and its intrinsics and extrinsics are set, we render the 3D occupancy into the corresponding view. Our goal is fast, fully differentiable rendering that supports efficient gradient backpropagation. To meet this need, we adopt a Gaussian splatting approach [23], which is significantly faster than traditional ray-casting while preserving differentiability. Consequently, we represent each voxel as a Gaussian primitive.
@@
To derive a Gaussian from a voxel, we emphasize simplicity to avoid degenerate configurations and ease the learning process. In practice, we: (1) use spherical Gaussians: represent each voxel as a sphere, which removes the need for orientation parameters; (2) fix the center: place each Gaussian at the center of its corresponding voxel, eliminating the need to learn an offset; (3) fix the scale: set the scale of each Gaussian based on the voxel dimensions.
@@ 4.1 Data and models
Occ3D-nuScenes is also based on the nuScenes dataset. It contains 18 semantic classes and has a 40cm voxel grid of range [-$[\text{-}40,40]\!\times\![\text{-}40,40]\!\times\![\text{-}1,5.4]$ meters. One major difference with SurroundOcc-nuScenes, is that Occ3D only evaluates the voxels visible from the cameras at the current time frame. Thus, it focuses on the geometric and semantic understanding of the visible objects, rather than extrapolating to occluded regions, leading to a simpler task.
SSCBench-Kitti360 [31] is derived from the Kitti360 dataset [34], acquired in Germany. It contains 19 semantic classes and has a 20cm voxel grid of range $[0,51.2]\!\times\![-25.6,25.6]\!\times\![-2,4.4]$ meters, resulting in a very precise semantic of urban scenes. It evaluates both visible and occluded voxels.
@@
We evaluate GaussRender across multiple models [54, 15, 19] and datasets [54, 50, 31] for 3D semantic occupancy prediction. Our method consistently improves performance without relying on other sensors such as the lidar used in certain methods [43, 49]. Results are summarized in Tab. 1 and detailed scores by class can be found in App. E.
@@
SurroundOcc-nuScenes [54]. On this dataset, considering visible and occluded voxels, GaussRender brings significant gains to all tested models. As shown in Tab. 1, TPVFormer [15] and SurroundOcc [54] reach the top two ranks. They achieve higher IoU (+1.2% and +1.1%) and mIoU (+3.8% and +0.5%) compared to their original implementations. Remarkably, GaussRender enables these models to surpass more recent approaches like GaussianFormerV2 [16] in both IoU and mIoU, proving that older architectures can achieve state-of-the-art results when their training is enhanced with GaussRender.
Models                       Frames RayIoU RayIoU1m, 2m, 4m
RenderOcc [43] 1 19.5 13.4, 19.6, 25.5
BEVDet-Occ [13] 2 29.6 23.6, 30.0, 35.1
BEVFormer [32] 4 32.4 26.1, 32.9, 38.0
BEVDet-Occ-Long [13] 8 32.6 26.6, 33.1, 38.2
SparseOcc[36] 8 34.0 28.0, 34.7, 39.4
FB-Occ [33] 16 33.5 26.7, 34.1, 39.7
SparseOcc [36] 16 36.1 30.2, 36.8, 41.2
TPVFormer [15] 1 37.2 31.5, 38.1, 41.9
w/ GaussRender 1 38.3 +1.1 32.3, 39.3, 43.4
SurroundOcc [54] 1 35.5 29.9, 36.3, 40.1
w/ GaussRender 1 37.5 +2.0 31.4, 38.5, 42.6
Table 2: Impact of GaussRender on RayIoU ($\uparrow$) metrics on the Occ3d-nuscenes validation dataset. Best results are in bold. Previous results are reported from [36]. ‘Frames’ indicates the number of frames used by the method. RayIoU1m, 2m, 4m refers to the RayIoU with a depth tolerance of 1 meter, 2 meters or 4 meters.
@@
Occ3D-nuScenes. On Occ3D-nuScenes [50], filtering unobserved voxels, GaussRender leads to the top two results: TPVFormer with GaussRender achieves 30.48 mIoU (+2.65), ranking first, and SurroundOcc with GaussRender reaches 30.38 mIoU (+1.17), ranking second. Notably, our approach outperforms RenderOcc [43] and GSRender [49] without requiring lidar inputs for loss computation. This demonstrates that methods augmented with GaussRender can learn strong 3D representations using only cameras.
@@
SSCBench-Kitti360. On the challenging SSCBench-Kitti360 [34] dataset, integrating GaussRender to SurroundOcc [54] and Symphonies [19] boosts IoU (+0.11% and +0.68%) and mIoU (+0.26% and +0.29%). While absolute gains appear smaller here, models in this benchmark achieve tightly clustered performance due to the smaller voxel size making the segmentation task more difficult; even modest improvements are difficult to achieve.

### v3

@@ Abstract
Understanding the 3D geometry and semantics of driving scenes is critical for safe autonomous driving. Recent advances in 3D occupancy prediction have improved scene representation but often suffer from visual inconsistencies, leading to floating artifacts and poor surface localization. Existing voxel-wise losses (e.g., cross-entropy) fail to enforce visible geometric coherence. In this paper, we propose GaussRender, a module that improves 3D occupancy learning by enforcing projective consistency. Our key idea is to project both predicted and ground-truth 3D occupancy into 2D camera views, where we apply supervision. Our method penalizes 3D configurations that produce inconsistent 2D projections, thereby enforcing a more coherent 3D structure. To achieve this efficiently, we leverage differentiable rendering with Gaussian splatting. GaussRender seamlessly integrates with existing architectures while maintaining efficiency and requiring no inference-time modifications. Extensive evaluations on multiple benchmarks (SurroundOcc-nuScenes, Occ3D-nuScenes, SSCBench-KITTI360) demonstrate that GaussRender significantly improves geometric fidelity across various 3D occupancy models (TPVFormer, SurroundOcc, Symphonies), achieving state-of-the-art results, particularly on surface-sensitive metrics such as RayIoU. The code is open-sourced at https://github.com/valeoai/GaussRender.
Figure 1: Comparison of rendered 3D predictions. Standard 3D Occupancy models trained on per-voxel losses result in physically implausible predictions (e.g., floating voxels, poorly localized surfaces, highlighted with orange ellipses) that maintain high 3D IoU but fail to produce visually consistent predictions. GaussRender enforces multi-view consistency, eliminating artifacts through learning projective constraints.
@@ 1 Introduction
The challenge in 3D occupancy prediction lies in achieving geometrically coherent reasoning from multi-view images. Existing methods [54, 15, 16, 30] typically optimize per-voxel predictions using standard 3D losses, such as cross-entropy, Dice [48], or Lovász [2]. However, these losses treat all voxels equally and do not enforce spatial consistency between neighboring voxels, leading to visual artifacts, as noticed in prior works [36] and illustrated in Fig. 1. Such artifacts (floating voxels, disjoint surfaces, and misaligned boundaries) have a low impact on voxel-based segmentation losses, which give the same weight to all voxels. The consequences extend beyond mere visual artifacts: poor surface localization and unrealistic floating objects may significantly undermine downstream tasks such as free-space estimation or motion planning. The underlying issue is that conventional supervision methods lack mechanisms to penalize unrealistic spatial arrangements, a deficiency that becomes especially apparent when projecting 3D artifacts into 2D views. This observation motivates our key insight: integrating projective consistency into the training objective encourages the model to learn consistent, physically and visually plausible geometries.
@@
We introduce GaussRender, a module that bridges the gap through differentiable rendering of 3D occupancy predictions. Our key idea is to enforce spatial consistency by projecting both predicted and ground-truth voxels into 2D camera views using Gaussian splatting [23]. These projections serve, during training, as a supervision signal, allowing us to penalize 3D configurations that produce poor 2D projections, thereby enforcing a more coherent and geometrically plausible 3D structure. To achieve this, we apply two complementary supervision signals: (1) a semantic rendering loss that enforces local semantic coherence, and (2) a depth rendering loss that penalizes occlusion-disrupting artifacts. These rendering-based losses are applied alongside standard 3D supervision, such as cross-entropy, ensuring compatibility with existing occupancy learning frameworks. GaussRender enables rendering from arbitrary viewpoints. Its flexibility helps reduce occlusions, for instance, by leveraging elevated viewpoints that are less affected by horizontal obstructions from ground objects.
@@
We validate our approach on multiple datasets, including SurroundOcc-nuScenes [54], Occ3D-nuScenes [50], and SSCBench-KITTI360 [31]. Our experiments demonstrate that GaussRender significantly improves geometric fidelity across diverse architectures — ranging from multi-scale voxel-based models (SurroundOcc [54]), tri-plane models (TPVFormer [15]) to hybrid query-voxel models (Symphonies [19]). Across all settings and datasets, GaussRender consistently improves performance on classical metrics such as IoU and mIoU. Moreover, when evaluated using surface- and artifact-sensitive metrics, such as RayIoU [36], the improvements are even more pronounced. This is because the projective constraints enforced by our method promote surface continuity — ensuring that neighboring voxels agree when rendered from any viewpoint, thereby eliminating floating artifacts and discontinuities. Crucially, our approach is plug-and-play, integrating seamlessly with existing 3D occupancy frameworks without requiring any architectural modifications. Thanks to our efficient Gaussian rendering proxy, GaussRender incurs minimal computational and memory overhead compared to NeRF-based alternatives [17, 43].
@@
  • A rendering-based module that enforces semantic and geometric consistency in 3D occupancy prediction, eliminating visual and spatial artifacts through projective constraints,
@@
  • A camera positioning strategy that amplifies supervision signals in geometrically complex regions.
  • Improves results on three standard benchmarks for many models, with particular gains in RayIoU.
@@ 2.1 Learning 3D semantic geometry from cameras
Discrete methods rely on regular 3D grids to stay as close as possible to the desired voxel output. Bird’s-eye-view (BeV) projections [51, 27] significantly reduce memory cost but introduce a substantial bias in the intermediate representation due to height compression. Instead, tri-plane features [15] and tensor decompositions [62] map any 3D coordinate onto distinct planes or vectors performing interpolation to get distinct features. As opposed to BeV approaches, they do not suffer from height compression while still operating on a compressed representation. Octrees [39] preserves a full multi-scale 3D representation, coarse in uniform areas and fine where details are needed, resulting in a fast and memory-efficient representation [54]. Continuous architectures have emerged, using a set of Gaussians to represent the scene [16, 18] which are then discretized to obtain the voxelized output map. Furthermore, to better capture the overall instance semantics and scene context, query-based methods have been introduced [19]. They facilitate interactions between image and volume features using sparse instance queries.
@@
Figure 2: Overview of GaussRender. Our module enforces 3D-2D consistency via differentiable Gaussian rendering. First, both predicted and ground-truth voxel grids are ‘gaussianized’ by converting each voxel into a simple spherical Gaussian: the center $\mu$ is fixed at the voxel center, the scale $s$ is a simple fixed scaling of the original voxel dimensions, and features are directly transferred as the semantic ‘classes’ $c$ — with only the opacity $o$ learned when voxel features are available. Next, virtual cameras are positioned in the scene (a fixed bird’s-eye view and a dynamic, arbitrarily placed camera as described in Sec. 3.2). The resulting 3D Gaussians are then projected into 2D using Gaussian splatting (Sec. 3.3), producing both semantic and depth renderings. These rendered views are compared against their ground-truth counterparts using an L1 penalty, ensuring enhanced spatial coherence and geometric consistency (Sec. 3.4).
@@
Complementary approaches have been proposed to enhance 3D occupancy predictions beyond architectural choices. Temporal aggregation uses past frames to resolve occlusions and refine geometric details [27, 59, 26, 46, 43], with extensions to 4D forecasting for dynamic scenes [24, 58, 53, 40]. Self-supervised methods reduces dependency on 3D annotations by generating pseudo-labels from monocular depth and segmentation [17, 7], but suffer from scale miscalibrations and label mismatches between 2D and 3D labels. Lastly, supervised rendering with lidar reprojections [43, 49] has been introduced to provide accurate depths. While more precise than pseudo-annotations, lidar reprojections face challenges such as signal sparsity, occlusions, and misalignment between the lidar and cameras. In particular, [43, 49] require temporal supervision from adjacent frames to compensate for lidar sparsity and occlusions. In contrast, our method achieves accurate supervision without these constraints, allowing more flexible and efficient 3D occupancy learning across different architectures.
@@ 3.1 Enforcing 3D-2D consistency
While the standard 3D losses (denoted as $L_{\text{3D}}$) ensure overall occupancy alignment with the ground truth, they treat each voxel independently and ignore spatial consistency. This limitation can lead to artifacts such as floating voxels or misaligned surfaces [36]. To overcome these issues, we introduce a 2D rendering loss, $L_{\text{2D}}$, which provides additional supervision by comparing rendered views of the predicted occupancy with corresponding ground-truth images. This loss enforces spatial coherence in the 2D projection, thereby penalizing 3D configurations that produce poor renderings. The overall training loss is defined as:
@@ 3.2 Camera Placement Strategy
GaussRender is not restricted to the original sensor or adjacent frames [43, 17, 49], it renders images from arbitrary camera positions within the scene. This flexibility allows us to position virtual cameras anywhere in the 3D space. Fig. 3 illustrates how virtual cameras generate complementary constraints (both semantic and depth) to enhance voxel consistency.
The placement of these virtual cameras is crucial and has been extensively studied. Several strategies have been found depending on the objective:
  • Visible Voxels: If the goal is to accurately reconstruct only the visible portions of the scene [50, 31], the rendering cameras should be placed close to the original sensor positions, as illustrated in Fig. 3 (a) in order to constraint the visibile voxels. This placement ensures that the inferred 3D structure remains consistent with the sensor’s view.
  • Holistic 3D Reconstruction: For complete 3D reconstruction, including occluded regions [54], cameras should be positioned more diversely. Exploring different viewpoints provides additional supervision for occluded areas, heavily penalizing aberrant configurations. In practice, we generate the virtual views with a simple rule: the camera is (1) elevated along $z$ axis with respect to its original position, (2) randomly translated in a close range to the ego-vehicle on the $xy$ plane. The resulting camera increases the field of view, looking at both visible and occluded parts of the scene (Fig. 3 (b)).
@@ 3.3 Gaussian rendering
For each camera, once its position is chosen and its intrinsics and extrinsics are set, we render the 3D occupancy into the corresponding view. Our goal is fast, fully differentiable rendering that supports efficient gradient backpropagation. To do it, we adopt a Gaussian splatting approach [23], which is significantly faster than traditional ray-casting while preserving differentiability. Consequently, we represent each voxel as a Gaussian primitive.
@@
To derive a Gaussian from a voxel, we emphasize simplicity to ease the learning process. In practice, we: (1) use spherical Gaussians: represent each voxel as a sphere, which removes the need for orientation parameters; (2) fix the center: place each Gaussian at the center of its corresponding voxel, eliminating the need to learn an offset; (3) fix the scale: set the scale of each Gaussian based on the voxel dimensions.
@@ 4.1 Data and models
Occ3D-nuScenes is also based on the nuScenes dataset. It contains 18 semantic classes and has a 40cm voxel grid of range [-$[\text{-}40,40]\!\times\![\text{-}40,40]\!\times\![\text{-}1,5.4]$ meters. One major difference with SurroundOcc-nuScenes, is that it only evaluates the voxels visible from the cameras at the current time frame. Thus, it focuses on the geometric and semantic understanding of the visible objects, rather than extrapolating to occluded regions, leading to a simpler task.
SSCBench-Kitti360 [31] is derived from the Kitti360 dataset [34], acquired in Germany. It contains 19 semantic classes and has a 20cm voxel grid of range $[0,51.2]\!\times\![-25.6,25.6]\!\times\![-2,4.4]$ meters, resulting in a very precise semantic of urban scenes. It evaluates both visible and occluded voxels, making the dataset particularly challenging due to its voxel resolution and the presence of occlusions.
@@
We evaluate GaussRender across multiple models [54, 15, 19] and datasets [54, 50, 31] for 3D semantic occupancy prediction. Our method consistently improves performance without relying on other sensors such as the lidar used in other works [43, 49]. Results are summarized in Tab. 1 and detailed scores by class can be found in App. F.
@@
SurroundOcc-nuScenes [54]. On this dataset, considering visible and occluded voxels, GaussRender brings significant gains to all tested models. As shown in Tab. 1, TPVFormer [15] and SurroundOcc [54] reach the top two ranks. They achieve higher IoU (+1.2 and +1.1) and mIoU (+3.8 and +0.5) compared to their original implementations. Remarkably, GaussRender enables these models to surpass more recent approaches like GaussianFormer and GaussianFormerV2 [16] in both IoU and mIoU, proving that older architectures can achieve state-of-the-art results when their training is enhanced with GaussRender.
@@
Occ3D-nuScenes. On Occ3D-nuScenes [50], GaussRender leads to the top two results: TPVFormer with GaussRender achieves 30.48 mIoU (+2.65), ranking first, and SurroundOcc with GaussRender reaches 30.38 mIoU (+1.17), ranking second. Notably, our approach outperforms other rendering methods such as RenderOcc [43] and GSRender [49] without requiring lidar inputs for loss computation. In addition, we also evaluate our model using only the rendering losses and compare to other supervised and unsupervised methods, seeTab. 2. While being static (single frame), it outperforms both static (+6.0 mIoU) and temporal (+1.4 mIoU) versions of RenderOcc, the latter casting rays from other frames achieving a new state-of-the art score on 2D supervision for 3D Occupancy on Occ3d-nuScenes. This demonstrates that methods augmented with GaussRender can learn strong 3D representations using only cameras.
Methods Labels 2D GT Img. Train Mem. Overhead Single Frame mIoU ($\uparrow$)
SelfOcc Pseudo-lbs. Dense High 9.3
OccNerf Pseudo-lbs. Dense High 9.5
GaussianOcc Pseudo-lbs. Dense Low 9.9
RenderOcc Lidar Sparse High 19.3 (S) / 23.9 (T)
GaussRender Voxels Dense Low 25.3
Table 2: Comparison of rendering-based methods on Occ3d-nuScenes under 2D-only supervision. We report the mIoU on fully trained models. Labels include pseudo-labels, sparse Lidar, or dense voxel supervision. GaussRender achieves the best performance with low memory and using a dense image supervision signal.
@@
SSCBench-Kitti360. On the challenging SSCBench-Kitti360 [34] dataset, integrating GaussRender to SurroundOcc [54] and Symphonies [19] boosts IoU (+0.11 and +0.68) and mIoU (+0.26 and +0.29). Absolute gains are smaller because models in this benchmark achieve tightly clustered performance due to the smaller voxel size making the segmentation task more difficult. However, when we look only at the visible voxels or in the sensor image metrics, we better see the overall improvements as figured in Tab. 3. SurroundOcc with GaussRender achieves higher mIoU (+1.23) and IoU (+1.99) on visible voxels. It clearly appears that GaussRender has strong impact on visible voxels meaning it has removed visibility artefacts.
3D all mIoU / IoU 3D visible mIoU / IoU Image mIoU / IoU Depth L1
SurroundOcc [54] 13.08 / 38.51 25.83 / 69.40 34.20 / 90.72 1.394
w/ GaussRender 13.34 / 38.62 27.06 / 71.39 35.86 / 94.55 1.173
+0.26 / +0.11 +1.23 / +1.99 +1.66 / +3.83 -0.221
Table 3: Performance improvement of SurroundOcc using GaussRender on KITTI-360. GaussRender leads to clear gains even on challenging datasets, especially in visible 3D regions.